# Minimax estimation of functionals in sparse vector model with correlated observations

### v1

@@ 3 Estimating the noise level
Thus, the minimum of $\sigma_{i}^{2}$’s is controlled by a positive constant that does not depend on $d$. It implies that $\|{{\boldsymbol{\Sigma}}}\|_{F}\gtrsim\sqrt{d}$. On the other hand, in this section, we do not need Assumption $(\varepsilon^{*}0)$.
Under Assumption $(\varepsilon^{*}1)$, it is convenient to work with the normalized observations $\tilde{Y}_{i}=\tilde{\theta}_{i}+\sigma\tilde{\varepsilon}_{i}$ for $i\in[d]$, where $\tilde{\theta}_{i}=\theta_{i}/\sigma_{i}$, and $\tilde{\varepsilon}_{i}=\varepsilon_{i}/\sigma_{i}$ is a standard Gaussian random variable. Thus, we replace the covariance matrix ${{\boldsymbol{\Sigma}}}$ by the correlation matrix $\tilde{{\boldsymbol{\Sigma}}}$ with entries
$\tilde{{\boldsymbol{\Sigma}}}_{ij}={\boldsymbol{\Sigma}}_{ij}/(\sigma_{i}\sigma_{j}),\quad i,j\in[d].$
@@ 3.1 Estimating $\sigma^{2}$ in the sparse regime
A relevant choice of $t$ in (11) is to be determined. Notice that for any deterministic $t>0$ this estimator is not well-defined since the denominator in (11) can be 0 with positive probability. Indeed, for any $t>0$ there exist ${\boldsymbol{\theta}}\in B_{0}(s,d)$ such that ${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}[\hat{F}_{Y^{2}}(t)=0]>0$. Thus, we will only define $\hat{\sigma}_{S}^{2}(t)$ by (11) if the denominator in (11) is non-zero and we set $\hat{\sigma}_{S}^{2}(t)=+\infty$ otherwise. The next lemma provides a bound on the estimation error of $\hat{\sigma}_{S}^{2}(t)$ for any fixed $t>0$.
@@
There exists an absolute constant $C>0$ such that for any ${\boldsymbol{\theta}}\in B_{0}(s,d)$ and any ${\boldsymbol{\Sigma}}$ satisfying Assumption LABEL:as:var1 the following holds.
(i) For any $t>0$ and $20\leq u\leq 1/$ 2 we have
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\left|\frac{\hat{\sigma}_{S}^{2}(t)}{\sigma^{2}}-1\right|\geq u\right]\leq\frac{C}{u^{2}}\left(\frac{s^{2}}{d^{2}}+\frac{\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}^{2}((\frac{t}{\sigma^{2}})^{2}+1)e^{-\frac{t}{3\sigma^{2}}}}{d^{2}}\right)\frac{e^{2t/\sigma^{2}}}{t/\sigma^{2}}.$
@@
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\left|\frac{\hat{\sigma}_{S}^{2}(t)}{\sigma^{2}}-1\right|\geq\frac{1}{2}\right]\leq C\frac{\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}^{2}}{d^{2}}.$
Lemma 1 suggests that the estimator $\hat{\sigma}_{S}^{2}(t)$ achieves a reasonable accuracy for $t\asymp\sigma^{2}$. However, such choice of $t$ is not realizable since $\sigma$ is unknown. Moreover, for any deterministic $t>0$ the expected error of $\hat{\sigma}_{S}^{2}(t)$ is not bounded uniformly over $\boldsymbol{\theta}$ since $\hat{\sigma}_{S}^{2}(t)$ can be +$\infty$+ ∞ with some positive probability as pointed out above. Therefore, we turn to an alternative solution that consists in choosing a random $t=\hat{t}$ based on the sample. Note that the standard sample splitting strategy, under which $\hat{t}$ can be independently chosen, is not a viable option since we have a model with dependent observations. We avoid sample splitting and propose a data-driven threshold $\hat{t}$ determined by scanning $\hat{F}_{Y^{2}}$ over a dyadic grid, where $\hat{F}_{Y^{2}}$ is based on the same set of observations as $\hat{\sigma}_{S}^{2}(\cdot)$ 111 The choice of constant 0.5 in the definition of $\hat{t}$ is not crucial. Replacing 0.5 by any other constant within $(0,1)$ only changes the resulting rates by a constant factor. :
@@
There exist absolute constants $C>0$ and $c>0$ such that for any $s\leq cd$, any ${\boldsymbol{\theta}}\in B_{0}(s,d)$ and any ${\boldsymbol{\Sigma}}$ satisfying Assumption LABEL:as:var1 we have
${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\frac{|\hat{\sigma}_{S}^{2}-\sigma^{2}|}{\sigma^{2}}\right]\leq C\left(\frac{s}{d}+\frac{\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}}{d}\right).$
It can be seen from the proof of Theorem 5 that the threshold (12) obtained by a search on a dyadic grid satisfies $\hat{t}\asymp\sigma^{2}$ with high probability. An alternative approach could be to use the sample median $\hat{t}=\min\{t:\hat{F}_{Y^{2}}(t)\geq 0.5\}$, which is also of the order of $\sigma^{2}$ with high probability. The main limitation of this approach is that, under the dependent data, the resulting estimator $\hat{\sigma}^{2}(\hat{t})$ becomes difficult to analyze. On the other hand, as we prove below, the value $\hat{t}$ defined in (12) belongs with high probability to a finite set
$\mathcal{T}:=\left\{t=2^{\ell}:\ell\in\mathbb{Z},\frac{\sigma^{2}}{3}\leq t\leq\frac{3}{2}\sigma^{2}\right\}.$
This allows us to control ${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\frac{|\hat{\sigma}_{S}^{2}-\sigma^{2}|}{\sigma^{2}}\right]$ by analyzing the convergence of $\hat{F}_{Y^{2}}(t)$ for every fixed $t\in\mathcal{T}$. In other words, with our dyadic grid search we restrict the “degrees of freedom” in choosing $\hat{t}$, which makes the resulting estimator more robust against data dependence.
It follows from Theorem 5 that in the sparse regime $s\leq\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}$ the risk of our the estimator $\hat{\sigma}_{S}^{2}$ is of order $\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}/d$, which is the optimal rate in this regime (see Corollary 2 below). However, in the dense regime $s\geq\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}$, the risk of $\hat{\sigma}_{S}^{2}$ scales as $s/d$, which is not optimal. In the next section, we provide another estimator of $\sigma^{2}$ that outperforms $\hat{\sigma}_{S}^{2}$ in the dense regime and attains the minimax optimal rate in both regimes.
3.2 Optimal estimation of the noise level
@@
$\hat{\varphi}_{d}(u):=\frac{1}{d}\sum_{i=1}^{d}\cos(u\tilde{Y}_{i}),\quad u>0.$
Analogously to (10), when $s$ is small compared to $d$ we can approximate $|\hat{\varphi}_{d}(u)|$ by $|\frac{1}{d}\sum_{i=1}^{d}\cos(u\sigma\tilde{\varepsilon}_{i})|$, which in turn is approximated by the value of the normal characteristic function $\varphi(u)=\exp(-u^{2}\sigma^{2}/2)$. Solving $\exp(-u^{2}\sigma^{2}/2)\approx|\hat{\varphi}_{d}(u)|$ suggests $-\frac{2}{u^{2}}\log|\hat{\varphi}_{d}(u)|$ as an estimator of $\sigma^{2}$, under some choice of $u$.
@@
$\tilde{\sigma}^{2}_{D}:=-\frac{2\hat{t}}{\lambda}\log\Big{|}\hat{\varphi}_{d}\Big{(}\sqrt{\lambda/\hat{t}}\Big{)}\Big{|},$
@@
There exist absolute constants $C>0$ and $c>0$ such that for any $s\leq cd$, any ${\boldsymbol{\theta}}\in B_{0}(s,d)$ and any ${\boldsymbol{\Sigma}}$ satisfying Assumption LABEL:as:var1 we have
${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\frac{|\hat{\sigma}_{D}^{2}-\sigma^{2}|}{\sigma^{2}}\right]\leq C\left(\frac{s}{d(1\vee\log({s}/{\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}}))}+\frac{\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}}{d}\right).$
@@
$\tilde{\psi}(s,{\boldsymbol{\Sigma}}):=\left\{\begin{array}[]{lc}\frac{\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}}{d}&\textrm{if}\;s\leq\|\tilde{{\boldsymbol{\Sigma}}}\|_{F},\\ \frac{s}{d(1\vee\log(s/\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}))}&\textrm{otherwise.}\end{array}\right.$
The following corollary shows that $\tilde{\psi}(s,{\boldsymbol{\Sigma}})$ is the minimax rate on the class $\Theta(s,d)$.
Corollary 2.
We have
$\inf_{\hat{T}}\sup_{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in\Theta(s,d)}{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\tilde{\psi}(s,{\boldsymbol{\Sigma}})^{-1}\frac{|\hat{\sigma}_{D}^{2}-\sigma^{2}|}{\sigma^{2}}\right]\asymp 1.$
This corollary follows immediately from Theorem 6 and the lower bound proved in  [8, Proposition 7] for $\tilde{{\boldsymbol{\Sigma}}}={\boldsymbol{\Sigma}}={\boldsymbol{I}}$.
@@ 4 Estimating the $\ell_{2}$-norm when the noise level is unknown
$\psi^{*}(s,{\boldsymbol{\Sigma}}):=\phi^{*}(s,\|{\boldsymbol{\Sigma}}\|_{F}^{2})=\left\{\begin{array}[]{lc}s\log(1+\|{\boldsymbol{\Sigma}}\|_{F}^{2}/s^{2})&\textrm{if}\;s\leq\|{\boldsymbol{\Sigma}}\|_{F},\\ \frac{s}{1\vee\log(s/\|{\boldsymbol{\Sigma}}\|_{F})}&\textrm{otherwise.}\end{array}\right.$
Our aim in this section is to show that $\sigma^{2}\psi^{*}(s,{\boldsymbol{\Sigma}})$ is the minimax optimal rate for the squared error of estimating $\|{\boldsymbol{\theta}}\|_{2}$ on $\Theta(s,d)$ if Assumptions LABEL:as:var0 and LABEL:as:var1 hold, and that the estimator $\hat{N}^{*}$ or its modifications attain this rate. The following theorem provides bounds on the accuracy of the estimator $\hat{N}^{*}$.
@@
There exist an absolute constant $\alpha\in(0,1)$ such that, for any ${\boldsymbol{\theta}}\in B_{0}(s,d)$ with $s$ satisfying that $s\leq\alpha d$ and any ${\boldsymbol{\Sigma}}$ satisfying Assumptions LABEL:as:var0 and LABEL:as:var1, the following holds.
@@
    ${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}[(\hat{N}^{*}-\|{\boldsymbol{\theta}}\|_{2})^{2}]\leq c_{1}\sigma^{2}\psi^{*}(s,{\boldsymbol{\Sigma}}).$
@@
$\hat{\sigma}_{\eta}^{2}:=\frac{\hat{M}}{q_{1-\eta/20}},\quad\textrm{where}\quad\hat{M}:=\min\{t:\hat{F}_{Y}^{2}(t)\geq 0.5\}.$
@@
$\hat{Q}^{*}_{\eta}:=\sum_{i=1}^{d}Y_{i}^{2}\mathbbm{1}\{|Y_{i}|>\max\{\hat{\sigma}_{S},\hat{\sigma}_{\eta}\}\sigma_{i}\tau_{\eta}\}-\hat{\sigma}_{S}^{2}\alpha_{s,\eta}\sum_{i=1}^{d}\sigma_{i}^{2},$
@@
$\tau_{\eta}=3\sqrt{\frac{q_{1-\eta/20}}{q_{\eta/20}}\log(1+\|{\boldsymbol{\Sigma}}\|_{F}^{2}/s^{2})}\quad\textrm{and}\quad\alpha_{s,\eta}={\mathbf{E}}[Z^{2}\mathbbm{1}\{|Z|\geq\tau_{\eta}\}].$
@@
There exist a constant $\alpha\in(0,1)$ such that for any $\eta\in(0,1)$, any ${\boldsymbol{\theta}}\in B_{0}(s,d)$ with $s$ satisfying that $s\leq\min\{\alpha d,\|{\boldsymbol{\Sigma}}\|_{F}\}$, and any ${\boldsymbol{\Sigma}}$ satisfying Assumptions LABEL:as:var0 and LABEL:as:var1 we have
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left((\hat{N}^{*}_{\eta}-\|{\boldsymbol{\theta}}\|_{2})^{2}\leq c_{\eta}\sigma^{2}\psi^{*}(s,{\boldsymbol{\Sigma}})\right)\geq 1-\eta,$
@@
We now discuss the lower bounds and minimax optimality. It follows from Theorem 7 (i) and Theorem 8 that the estimator $\hat{N}^{**}$, which equals to $\hat{N}^{*}$ if $s>\|{\boldsymbol{\Sigma}}\|_{F}$ and to $\hat{N}^{*}_{\eta}$ otherwise, satisfies
$\sup_{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in\Theta(s,d)}\sup_{\sigma>0}{\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\big{(}(\hat{N}^{**}-\|{\boldsymbol{\theta}}\|_{2})^{2}\geq C_{\eta}^{\prime}\sigma^{2}\psi^{*}(s,{\boldsymbol{\Sigma}})\big{)}\leq\eta$ (16)
for a constant $C_{\eta}^{\prime}>0$ depending only on $\eta$. On the other hand, we have a matching lower bound that follows immediately from Proposition 3 (ii) in [8] proved for the i.i.d. case (${\boldsymbol{\Sigma}}={\boldsymbol{I}}$): There exist constants $\eta^{*}\in(0,1)$, $c^{\prime}>0$ such that
$\inf_{\hat{T}}\sup_{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in\Theta(s,d)}\sup_{\sigma>0}{\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\big{(}(\hat{T}-\|{\boldsymbol{\theta}}\|_{2})^{2}\geq c^{\prime}\sigma^{2}\psi^{*}(s,{\boldsymbol{\Sigma}})\big{)}>\eta^{*},$ (17)
where $\inf_{\hat{T}}$ denotes the infimum over all estimators. Thus, the rate $\sigma^{2}\psi^{*}(s,{\boldsymbol{\Sigma}})$ is minimax optimal on $\Theta(s,d)$ in the sense that, at any confidence level $1-\eta$ (uniform over the class), where $\eta\in(0,\eta^{*})$, we can estimate $\|{\boldsymbol{\theta}}\|_{2}$ with an error not exceeding $C_{\eta}^{\prime}\sigma^{2}\psi^{*}(s,{\boldsymbol{\Sigma}})$ but we cannot estimate it with smaller error than $c^{\prime}\sigma^{2}\psi^{*}(s,{\boldsymbol{\Sigma}})$ at such confidence levels.
Next, we consider the setting where the Frobenius norm $\|{\boldsymbol{\Sigma}}\|_{F}$ is not available and we only know an upper bound $\rho$ on $\|{\boldsymbol{\Sigma}}\|_{F}$. In this setting, we deal with minimax optimality on the parameter spaces $\Theta(s,d,\rho)$, $\rho>0$, where the additional restriction $\|{\boldsymbol{\Sigma}}\|_{F}\leq\rho$ is active. Let $\tilde{N}^{*}$ and $\tilde{N}^{*}_{\eta}$ be the estimators defined in the same way as $\hat{N}^{*}$ and $\hat{N}^{*}_{\eta}$, respectively, with the only difference that $\|{\boldsymbol{\Sigma}}\|_{F}$ is replaced by its upper bound $\rho$. Let $\phi^{*}(s,\cdot)$ be defined in (3). Then $\phi^{*}(s,\rho^{2})$ follows the same definition as $\psi^{*}(s,{\boldsymbol{\Sigma}})$ except that we replace $\|{\boldsymbol{\Sigma}}\|_{F}$ by $\rho$.
@@
$\inf_{\hat{T}}\sup_{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in{\Theta}(s,d,\rho)}\sup_{\sigma>0}{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\,w\big{(}\sigma^{-1}(\phi^{*}(s,\rho^{2}))^{-1/2}|\hat{T}-\|{\boldsymbol{\theta}}\|_{2}|\big{)}\geq c.$
@@
    $\sup_{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in\Theta(s,d,\rho)}\sup_{\sigma>0}\frac{{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}[(\tilde{N}^{*}-\|{\boldsymbol{\theta}}\|_{2})^{2}]}{\sigma^{2}}\lesssim\phi^{*}(s,\rho^{2}).$
@@
    $\sup_{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in\Theta(s,d,\rho)}\sup_{\sigma>0}{\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left(({\tilde{N}^{*}_{\eta}}-\|{\boldsymbol{\theta}}\|_{2})^{2}\geq c_{\eta}\sigma^{2}\phi^{*}(s,\rho^{2})\right)\leq\eta.$

### v2

@@ 3 Estimating the noise level
Thus, the minimum of $\sigma_{i}^{2}$’s is controlled by a positive constant that does not depend on $d$. It implies that $\|{{\boldsymbol{\Sigma}}}\|_{F}\gtrsim\sqrt{d}$. On the other hand, in this section, we do not need Assumption 1.
Under Assumption 2, it is convenient to work with the normalized observations $\tilde{Y}_{i}=\tilde{\theta}_{i}+\sigma\tilde{\varepsilon}_{i}$ for $i\in[d]$, where $\tilde{\theta}_{i}=\theta_{i}/\sigma_{i}$, and $\tilde{\varepsilon}_{i}=\varepsilon_{i}/\sigma_{i}$ is a standard Gaussian random variable. Thus, we replace the covariance matrix $\boldsymbol{\Sigma}$}} by the correlation matrix $\tilde{{\boldsymbol{\Sigma}}}$ with entries
$\tilde{{\boldsymbol{\Sigma}}}_{ij}={\boldsymbol{\Sigma}}_{ij}/(\sigma_{i}\sigma_{j}),\quad i,j\in[d].$ (17)
@@ 3.1 Estimating $\sigma^{2}$ in the sparse regime
A relevant choice of $t$ in (19) is to be determined. Notice that for any deterministic $t>0$ this estimator is not well-defined since the denominator in (19) can be 0 with positive probability. Indeed, for any $t>0$ there exist ${\boldsymbol{\theta}}\in B_{0}(s,d)$ such that ${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}[\hat{F}_{Y^{2}}(t)=0]>0$. Thus, we will only define $\hat{\sigma}_{S}^{2}(t)$ by (19) if the denominator in (19) is non-zero. Otherwise, we set $\hat{\sigma}_{S}^{2}(t)=\sum_{i=1}^{d}Y_{i}^{2}/d$; we note that the conclusion remains valid for any other choice of this value. The next lemma provides a bound on the estimation error of $\hat{\sigma}_{S}^{2}(t)$ for any fixed $t>0$.
@@
There exists an absolute constant $C>0$ such that for any ${\boldsymbol{\theta}}\in B_{0}(s,d)$ and any $\boldsymbol{\Sigma}$} satisfying Assumption 2 the following holds.
(i) For any $t>0$ and $0\leq u\leq 1/2$ we have
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\left|\frac{\hat{\sigma}_{S}^{2}(t)}{\sigma^{2}}-1\right|\geq u\right]\leq\frac{C}{u^{2}}\left(\frac{s^{2}}{d^{2}}+\frac{\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}^{2}((\frac{t}{\sigma^{2}})^{2}+1)e^{-\frac{t}{3\sigma^{2}}}}{d^{2}}\right)\frac{e^{2t/\sigma^{2}}}{t/\sigma^{2}}.$ (20)
@@
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\left|\frac{\hat{\sigma}_{S}^{2}(t)}{\sigma^{2}}-1\right|\geq\frac{1}{2}\right]\leq C\frac{\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}^{2}}{d^{2}}.$ (21)
Lemma 1 suggests that the estimator $\hat{\sigma}_{S}^{2}(t)$ achieves a reasonable accuracy for $t\asymp\sigma^{2}$. However, such choice of $t$ is not realizable since $\sigma$ is unknown. Moreover, for any deterministic $t>0$ the expected error of $\hat{\sigma}_{S}^{2}(t)$ is not bounded uniformly over $\boldsymbol{\theta}$ since $\hat{\sigma}_{S}^{2}(t)$ can be $\infty$ with some positive probability as pointed out above. Therefore, we turn to an alternative solution that consists in choosing a random $t=\hat{t}$ based on the sample. Note that the standard sample splitting strategy, under which $\hat{t}$ can be independently chosen, is not a viable option since we have a model with dependent observations. We avoid sample splitting and propose a data-driven threshold $\hat{t}$ determined by scanning $\hat{F}_{Y^{2}}$ over a dyadic grid, where $\hat{F}_{Y^{2}}$ is based on the same set of observations as $\hat{\sigma}_{S}^{2}(\cdot)$  [^1] The choice of constant 0.5 in the definition of $\hat{t}$ is not crucial. Replacing 0.5 by any other constant within $(0,1)$ only changes the resulting rates by a constant factor. :
@@
There exist absolute constants $C>0$ and $c>0$ such that for any $s\leq cd$, any ${\boldsymbol{\theta}}\in B_{0}(s,d)$ and any $\boldsymbol{\Sigma}$} satisfying Assumption 2 we have
${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\frac{|\hat{\sigma}_{S}^{2}-\sigma^{2}|}{\sigma^{2}}\right]\leq C\left(\frac{s}{d}+\frac{\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}}{d}\right).$ (24)
[…]
$\mathcal{T}:=\left\{t=2^{\ell}:\ell\in\mathbb{Z},\frac{\sigma^{2}}{3}\leq t\leq\frac{3}{2}\sigma^{2}\right\}.$ (25)
This allows us to control ${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\frac{|\hat{\sigma}_{S}^{2}-\sigma^{2}|}{\sigma^{2}}\right]$ by analyzing the convergence of $\hat{F}_{Y^{2}}(t)$ for every fixed $t\in\mathcal{T}$. In other words, with our dyadic grid search we restrict the “degrees of freedom” in choosing $\hat{t}$, which makes the resulting estimator more robust against data dependence. In the next section, we provide another estimator of $\sigma^{2}$ that outperforms $\hat{\sigma}_{S}^{2}$ in the dense regime.
3.2 Estimating $\sigma^{2}$ in the dense regime
@@
$\hat{\varphi}_{d}(u):=\frac{1}{d}\sum_{i=1}^{d}\cos(u\tilde{Y}_{i}),\quad u>0.$ (26)
Analogously to (18), when $s$ is small compared to $d$ we can approximate $|\hat{\varphi}_{d}(u)|$ by $|\frac{1}{d}\sum_{i=1}^{d}\cos(u\sigma\tilde{\varepsilon}_{i})|$, which in turn is approximated by the value of the normal characteristic function $\varphi(u)=\exp(-u^{2}\sigma^{2}/2)$. Solving $\exp(-u^{2}\sigma^{2}/2)\approx|\hat{\varphi}_{d}(u)|$ suggests $\frac{2}{u^{2}}\log|\hat{\varphi}_{d}(u)|$ as an estimator of $\sigma^{2}$, under some choice of $u$.
@@
[…]
@@
There exist absolute constants $C>0$ and $c>0$ such that for any $s\leq cd$, any ${\boldsymbol{\theta}}\in B_{0}(s,d)$ and any $\boldsymbol{\Sigma}$} satisfying Assumption 2 we have
${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left[\frac{|\hat{\sigma}_{D}^{2}-\sigma^{2}|}{\sigma^{2}}\right]\leq C\left(\frac{s}{d(1\vee\log({s}/{\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}}))}+\frac{\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}}{d}\right).$ (28)
@@
$\tilde{\psi}(s,{\boldsymbol{\Sigma}}):=\left\{\begin{array}[]{lc}\frac{\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}}{d}&\textrm{if}\;s\leq\|\tilde{{\boldsymbol{\Sigma}}}\|_{F},\\ \frac{s}{d(1\vee\log(s/\|\tilde{{\boldsymbol{\Sigma}}}\|_{F}))}&\textrm{otherwise.}\end{array}\right$. (29)
@@ 4 Estimating the $\ell_{2}$-norm when the noise level is unknown
$\psi^{*}(s,{\boldsymbol{\Sigma}}):=\phi^{*}(s,\|{\boldsymbol{\Sigma}}\|_{F}^{2})=\left\{\begin{array}[]{lc}s\log(1+\|{\boldsymbol{\Sigma}}\|_{F}^{2}/s^{2})&\textrm{if}\;s\leq\|{\boldsymbol{\Sigma}}\|_{F},\\ \frac{s}{1\vee\log(s/\|{\boldsymbol{\Sigma}}\|_{F})}&\textrm{otherwise.}\end{array}\right$. (31)
Our aim in this section is to show that $\sigma^{2}\psi^{*}(s,{\boldsymbol{\Sigma}})$ is the minimax optimal rate for the squared error of estimating $\|{\boldsymbol{\theta}}\|_{2}$ on $\Theta(s,d)$ if Assumptions 1 and 2 hold, and that the estimator $\hat{N}^{*}$ or its modifications attain this rate. The following theorem provides bounds on the accuracy of the estimator $\hat{N}^{*}$.
@@
There exists an absolute constant $\alpha\in(0,1)$ such that, for any ${\boldsymbol{\theta}}\in B_{0}(s,d)$ with $s$ satisfying that $s\leq\alpha d$ and any $\boldsymbol{\Sigma}$} satisfying Assumptions 1 and 2, the following holds.
@@
    ${\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}[(\hat{N}^{*}-\|{\boldsymbol{\theta}}\|_{2})^{2}]\leq c_{1}\sigma^{2}\psi^{*}(s,{\boldsymbol{\Sigma}}).$ (32)
@@
$\hat{\sigma}_{\eta}^{2}:=\frac{\hat{M}}{q_{1-\eta/20}},\quad\textrm{where}\quad\hat{M}:=\min\{t:\hat{F}_{Y}^{2}(t)\geq 0.5\}.$ (35)
@@
$\hat{Q}^{*}_{\eta}:=\sum_{i=1}^{d}Y_{i}^{2}\mathbbm{1}\{|Y_{i}|>\max\{\hat{\sigma}_{S},\hat{\sigma}_{\eta}\}\sigma_{i}\tau_{\eta}\}-\hat{\sigma}_{S}^{2}\alpha_{s,\eta}\sum_{i=1}^{d}\sigma_{i}^{2},$ (36)
@@
$\tau_{\eta}=3\sqrt{\frac{q_{1-\eta/20}}{q_{\eta/20}}\log(1+\|{\boldsymbol{\Sigma}}\|_{F}^{2}/s^{2})}\quad\textrm{and}\quad\alpha_{s,\eta}={\mathbf{E}}[Z^{2}\mathbbm{1}\{|Z|\geq\tau_{\eta}\}].$ (37)
@@
There exists a constant $\alpha\in(0,1)$ such that for any $\eta\in(0,1)$, any ${\boldsymbol{\theta}}\in B_{0}(s,d)$ with $s$ satisfying that $s\leq\min\{\alpha d,\|{\boldsymbol{\Sigma}}\|_{F}\}$, and any $\boldsymbol{\Sigma}$} satisfying Assumptions 1 and 2 we have
${\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left((\hat{N}^{*}_{\eta}-\|{\boldsymbol{\theta}}\|_{2})^{2}\leq c_{\eta}\sigma^{2}\psi^{*}(s,{\boldsymbol{\Sigma}})\right)\geq 1-\eta,$ (38)
@@
Next, we consider the setting where the Frobenius norm $\|{\boldsymbol{\Sigma}}\|_{F}$ is not available and we only know an upper bound $\rho$ on $\|{\boldsymbol{\Sigma}}\|_{F}$. In this setting, we show that estimators that are based on slight modifications of $\hat{N}^{\star}$ and $\hat{N}^{\star}_{\eta}$ can achieve minimax optimality on the parameter spaces $\Theta(s,d,\rho)$, $\rho>0$, where the restriction $\|{\boldsymbol{\Sigma}}\|_{F}\leq\rho$ is active. Let $\tilde{N}^{*}$ and $\tilde{N}^{*}_{\eta}$ be the estimators defined in the same way as $\hat{N}^{*}$ and $\hat{N}^{*}_{\eta}$, respectively, with the only difference that $\|{\boldsymbol{\Sigma}}\|_{F}$ is replaced by its upper bound $\rho$. Let $\phi^{*}(s,\cdot)$ be defined in (5). Then $\phi^{*}(s,\rho^{2})$ follows the same definition as $\psi^{*}(s,{\boldsymbol{\Sigma}})$ except that we replace $\|{\boldsymbol{\Sigma}}\|_{F}$ by $\rho$.
@@
$\inf_{\hat{T}}\sup_{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in{\Theta}(s,d,\rho)}\sup_{\sigma>0}{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\,w\big(\sigma^{-1}(\phi^{*}(s,\rho^{2}))^{-1/2}|\hat{T}-\|{\boldsymbol{\theta}}\|_{2}|\big)\geq c.$ (39)
@@
    $\sup_{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in\Theta(s,d,\rho)}\sup_{\sigma>0}\frac{{\mathbf{E}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}[(\tilde{N}^{*}-\|{\boldsymbol{\theta}}\|_{2})^{2}]}{\sigma^{2}}\lesssim\phi^{*}(s,\rho^{2}).$ (40)
@@
    $\sup_{({\boldsymbol{\theta}},{\boldsymbol{\Sigma}})\in\Theta(s,d,\rho)}\sup_{\sigma>0}{\mathbf{P}}_{{\boldsymbol{\theta}},{\boldsymbol{\Sigma}}}\left(({\tilde{N}^{*}_{\eta}}-\|{\boldsymbol{\theta}}\|_{2})^{2}\geq c_{\eta}\sigma^{2}\phi^{*}(s,\rho^{2})\right)\leq\eta.$ (41)